Let f(x) = 1/x. Show that (f * f)(x) = x for all nonzero value of x.

Thank you for your help ahead of time.
It's very appeciated.

Answer by stanbon(48569)

Let f(x) = 1/x. Show that (f * f)(x) = x for all nonzero value of x.
--------------
Comment: (f*f)(x) means [f(x)*f(x)].
In your case that would be 1/x^2
I don't think you want that.
---------
I think what you want is fof(x) which means f[f(x)] = f(1/x)] 1/(1/x) = x
================
Cheers,
Stan H.